# Quantization Error Propagation: Revisiting Layer-Wise Post-Training Quantization

### v2

@@ A.2 Quantization Error Accumulation
We proceed by induction over the network layers.
Base Case
At layer $1$, the error clearly satisfies
$\|{\bm{\delta}}_{1}^{\mathrm{QEP}}\|_{F}=\|{\bm{\delta}}_{1}^{\mathrm{BASE}}\|_{F}=0,$ (52)
since $\widehat{{\bm{X}}}_{1}={\bm{X}}_{1}$ holds for both methods.
Inductive Step
Assume that for some layer $l\geq 1$
$\|{\bm{\delta}}_{l}^{\mathrm{QEP}}\|_{F}\leq\|{\bm{\delta}}_{l}^{\mathrm{BASE}}\|_{F}.$ (53)
Consider the error at layer $l+1$:
${\bm{\delta}}_{l+1}=\widehat{{\bm{X}}}_{l+1}-{\bm{X}}_{l+1}=\sigma_{l}(\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l})-\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l}).$ (54)
Using the Lipschitz continuity of $\sigma_{l}$, we bound the error as
$\|{\bm{\delta}}_{l+1}\|_{F}\leq\gamma_{l}\|\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}-{\bm{W}}_{l}{\bm{X}}_{l}\|_{F}.$ (55)
We analyze this bound separately for each quantization method. Let $\widehat{{\bm{X}}}^{\mathrm{BASE}}_{l}$ and $\widehat{{\bm{X}}}_{l}^{\mathrm{QEP}}$ denote the activations at layer $l$ for the BASE and QEP methods, respectively.
In BASE method, we have
$\widehat{{\bm{W}}}_{l}^{\mathrm{BASE}}\widehat{{\bm{X}}}_{l}^{\mathrm{BASE}}-{\bm{W}}_{l}{\bm{X}}_{l}$ $=({\bm{W}}_{l}+E_{l})\widehat{{\bm{X}}}_{l}^{\mathrm{BASE}}-{\bm{W}}_{l}{\bm{X}}_{l}$ (56)
$={\bm{W}}_{l}{\bm{\delta}}_{l}^{\mathrm{BASE}}+{\bm{E}}_{l}{\bm{X}}_{l}+{\mathcal{O}}({\bm{E}}_{l}{\bm{E}}_{l-1}).$ (57)
Including the correction term, we obtain
$\widehat{{\bm{W}}}_{l}^{\mathrm{QEP}}\widehat{{\bm{X}}}_{l}^{\mathrm{QEP}}-{\bm{W}}_{l}{\bm{X}}_{l}=\underbrace{{\bm{W}}_{l}{\bm{\delta}}_{l}^{\mathrm{QEP}}({\bm{I}}-\alpha_{l}{\bm{P}}_{l}^{\mathrm{QEP}})}_{{\bm{T}}_{l}^{(1)}}+{\bm{E}}_{l}{\bm{X}}_{l}+{\mathcal{O}}({\bm{E}}_{l}{\bm{E}}_{l-1}),$ (58)
where ${\bm{P}}_{l}^{\mathrm{QEP}}$ is a projection matrix, given by
$P_{l}^{\mathrm{QEP}}=\widehat{{\bm{X}}}_{l}^{\mathrm{QEP}\top}(\widehat{{\bm{X}}}_{l}^{\mathrm{QEP}}\widehat{{\bm{X}}}_{l}^{\mathrm{QEP\top}})^{-1}\widehat{{\bm{X}}}_{l}^{\mathrm{QEP}}.$ (59)
Regarding first term ${\bm{T}}^{(1)}_{l}$, because ${\bm{P}}_{l}$ is an orthogonal projection and ${\bm{0}}\leq\alpha_{l}\leq 1$, the eigenvalues of ${\bm{I}}-\alpha_{l}{\bm{P}}_{l}$ lie in $[0,1]$; therefore $\|{\bm{Z}}({\bm{I}}-\alpha_{l}{\bm{P}}\|_{F}\leq\|{\bm{Z}}\|$ for any ${\bm{Z}}$. Using this and the induction hypothesis,
$\|{\bm{T}}_{l}^{(1)}\|\leq\|{\bm{W}}_{l}\|_{2}\|{\bm{\delta}}^{\mathrm{QEP}}\|_{F}\leq\|{\bm{W}}_{l}\|_{2}\|{\bm{\delta}}^{\mathrm{BASE}}\|_{F}\leq\|{\bm{W}}_{l}{\bm{\delta}}_{l}^{B}\|_{F}.$ (60)
Regarding second term ${\bm{T}}_{l}^{(2)}$, apply the triangle inequality and the sub-multiplicativity $\|AB\|_{F}\leq\|A\|_{2}\|{\bm{B}}\|_{F}$:
$\|{\bm{T}}_{1}^{(2)}\|_{F}\leq\|{\bm{E}}_{l}{\bm{X}}_{l}\|_{F}+\|{\bm{E}}_{l}\|_{2}\|{\bm{\delta}}^{\mathrm{QEP}}\|_{F}.$ (61)
Thus, combining this property with the inductive hypothesis, we can derive
$\left\|{\bm{\delta}}_{l+1}^{\mathrm{QEP}}\right\|_{F}\leq\left\|{\bm{\delta}}_{l+1}^{\mathrm{BASE}}\right\|_{F}.$ (62)
By induction, the inequality holds for all layers, yielding
$\left\|f_{{\bm{\theta}}}({\bm{X}})-f_{\widehat{{\bm{\theta}}}^{\mathrm{QEP}}}({\bm{X}})\right\|_{F}\leq\left\|f_{{\bm{\theta}}}({\bm{X}})-f_{\widehat{{\bm{\theta}}}^{\mathrm{BASE}}}({\bm{X}})\right\|_{F}.$ (63)
This completes the proof. ∎
We further demonstrate that the final quantization error monotonically decreases as each propagation strength parameter $\alpha_{l}$ approaches $1$.
Corollary A.5.
Under the same assumptions as Theorem 5.2, consider two sets of QEP-corrected weights defined by parameters $\alpha_{l},\alpha_{l}^{\prime}$ with $0\leq\alpha_{l}^{\prime}\leq\alpha_{l}\leq 1$ for all layers $l=1,\dots,L$. We denote ${\bm{\alpha}}=(\alpha_{l})_{l=1}^{L}$. Then the following inequality holds:
$\|f_{{\bm{\theta}}}({\bm{X}})-f_{\widehat{{\bm{\theta}}}^{\mathrm{QEP}}_{{\bm{\alpha}}}}(X)\|_{F}\leq\|f_{{\bm{\theta}}}({\bm{X}})-f_{\widehat{{\bm{\theta}}}^{\mathrm{QEP}}_{{\bm{\alpha}}^{\prime}}}(X)\|_{F}.$ (64)
where ${\bm{\theta}}^{\mathrm{QEP}}_{{\bm{\alpha}}}$ represents QEP-quantized model parameters with ${\bm{\alpha}}$.
@@
The claim is proved by induction on the layer index $l$. For brevity the actication mismatch matrix ${\bm{\delta}}\coloneqq{\bm{X}}_{l}-\hat{{\bm{X}}}_{l}^{\alpha_{l}}$, $\hat{{\bm{X}}}_{l}^{\alpha_{l}}$ is the activation produced with the QEP-quantized weight that uses the parameter $\alpha_{l}$.
Base Case
At layer $1$, the error clearly satisfies
${\bm{\delta}}_{l}^{\alpha_{l}}={\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}^{\alpha_{l}},$ (65)
since ${\bm{\delta}}_{l}^{\alpha_{l}}={\bm{\delta}}_{l}^{\alpha_{l}^{\prime}}={\bm{X}}_{1}$ holds for both methods.
Inductive Step.
Suppose at layer $l\geq 1$ that
$\|{\bm{\delta}}_{l}^{\alpha_{l}}\|_{F}\leq\|{\bm{\delta}}_{l}^{\alpha_{l}^{\prime}}\|_{F}.$ (66)
Using the Lipschitz continuity of $\sigma_{l}(\cdot)$, we bound the error as
$\|{\bm{\delta}}_{l+1}^{\alpha_{l}}\|_{F}$ $\leq\gamma_{l}\|\widehat{{\bm{W}}}_{l}^{\alpha_{l}}\widehat{{\bm{X}}}_{l}^{\alpha_{l}}-{\bm{W}}_{l}{\bm{X}}_{l}\|_{F}.$ (67)
and an analogous bound with $\alpha_{i}^{\prime}$ in place of $\alpha_{l}$.
Split the right hand side:
$\widehat{{\bm{W}}}_{l}^{\alpha}\widehat{{\bm{X}}}_{l}^{\alpha}-{\bm{W}}_{l}{\bm{X}}_{l}={\bm{W}}_{l}{\bm{\delta}}_{l}^{\alpha_{l}}({\bm{I}}-\alpha_{l}{\bm{P}}_{l}^{\alpha_{l}})+{\bm{E}}_{l}{\bm{X}}_{l}+{\mathcal{O}}({\bm{E}}_{l}{\bm{E}}_{l-1}).$ (68)
Lemma A.7 states that for any matrix ${\bm{Z}}$ and $0\leq\alpha_{l}^{\prime}\leq\alpha_{l}\leq 1$,
$\|{\bm{Z}}({\bm{I}}-\alpha_{l}{\bm{P}}_{l}^{\alpha_{l}})\|_{F}\leq\|{\bm{Z}}({\bm{I}}-\alpha_{l}^{\prime}{\bm{P}}_{l}^{\alpha_{l}}\|_{F}.$ (69)
Applying this with ${\bm{Z}}={\bm{W}}_{l}{\bm{\delta}_{l}}^{\alpha_{l}}$ and using the induction hypothesis in Eq. (66) gives
$\|{\bm{W}}_{l}{\bm{\delta}}_{l}^{\alpha_{l}}({\bm{I}}-\alpha_{l}{\bm{P}}_{l}^{\alpha_{l}}\|_{F}\leq\|{\bm{W}}_{l}{\bm{\delta}}_{l}^{\alpha_{l}^{\prime}}({\bm{I}}-\alpha_{l}^{\prime}{\bm{P}}_{l}^{\alpha_{l}})\|_{F}.$ (70)
Combining this inequality and the inductive hypothesis, we obtain:
$\|\widehat{{\bm{W}}}_{l}^{\alpha}\widehat{{\bm{X}}}_{l}^{\alpha}-{\bm{W}}_{l}{\bm{X}}_{l}\|_{F}\leq\|\widehat{{\bm{W}}}_{l}^{\alpha^{\prime}}\widehat{{\bm{X}}}_{l}^{\alpha^{\prime}}-{\bm{W}}_{l}{\bm{X}}_{l}\|_{F},$ (71)
and consequently:
$\|{\bm{\delta}}_{l+1}^{\alpha_{l+1}}\|_{F}\leq\|{\bm{\delta}}_{l+1}^{\alpha^{\prime}_{l+1}}\|_{F}.$ (72)
Thus, by induction, we conclude that for every layer, increasing $\alpha_{l}$ closer to 1 monotonically reduces the activation mismatch and hence reduces the final quantization error. Formally,
$\left\|f_{{\bm{\theta}}}({\bm{X}})-f_{\widehat{\theta}^{\mathrm{QEP}}_{{\bm{\alpha}}}}({\bm{X}})\right\|_{F}\leq\left\|f_{\theta}({\bm{X}})-f_{\widehat{{\bm{\theta}}}^{\mathrm{QEP}}_{{\bm{\alpha}}^{\prime}}}({\bm{X}})\right\|_{F},$ (73)
as required. ∎
A.4 Relationship of QEP Correction and Ridge Regularization

### v3

@@ B.2 Quantization Error Accumulation
Since $\sigma_{l}$ is the identity map, we have
$X_{L}=(1+\varepsilon)^{L-1}C,~~\widehat{X}_{L}=(1+\varepsilon+c_{E})^{L-1}C.$ (62)
Hence
$|\delta_{L}|=C\left|(1+\varepsilon+c_{E})^{L-1}-(1+\varepsilon)^{L-1}\right|.$ (63)
Apply the mean value theorem to $f(t)=t^{L-1}$ on the interval $1+\varepsilon,\;1+\varepsilon+c_{E}]$: there exists $\xi\in(1+\varepsilon,1+\varepsilon+c_{E})$ such that
$(1+\varepsilon+c_{E})^{L-1}-(1+\varepsilon)^{L-1}=f^{\prime}(\xi)c_{E}=(L-1)\xi^{L-2}c_{E}.$ (64)
Since $\xi\geq 1+\varepsilon$, we obtain
$|\delta_{L}|\geq(L-1)c_{E}(1+\varepsilon)^{L-2}C,$ (65)
which proves Eq. (60). Finally, Eq. (61) follows from $(1+\varepsilon)^{L-2}=\nicefrac{{1}}{{1+\varepsilon}}(1+\varepsilon)^{L-1}$ and $L-1\geq 1$ for $L\geq 2$. ∎
B.3 Derivation of Theorem 5.2 and Corollary 5.4
This section presents a rigorous statement and proof of Theorem 5.2 and Corollary 5.4. We first formally restate Theorem 5.2 below.
Theorem B.5.
Consider an $L$-layer network
${\bm{X}}_{1}={\bm{X}},~~{\bm{X}}_{l+1}=\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l}),~~l=1,\dots,L.$ (66)
Assume each $\sigma_{l}$ is $\gamma_{l}$-Lipschitz with respect to $\|\cdot\|_{F}$ and satisfies $\sigma_{l}({\bm{0}})={\bm{0}}$. Let the quantized forward pass be
$\widehat{{\bm{X}}}_{1}={\bm{X}},~~\widehat{{\bm{X}}}_{l+1}=\sigma_{l}(\widehat{{\bm{W}}}_{l}\widehat{{\bm{X}}}_{l}).$ (67)
Define the activation mismatch as ${\bm{\delta}}_{l}\coloneq{\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$.
Fix any matrices ${\bm{E}}_{l}$ and set
$\widehat{{\bm{W}}}_{l}^{\mathrm{BASE}}\coloneq{\bm{W}}_{l}+{\bm{E}}_{l}.$ (68)
For QEP, define each $l$
${\bm{W}}_{l}^{\ast}(\alpha_{l})\coloneq{\bm{W}}_{l}+\alpha_{l}{\bm{W}}_{l}{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}\widehat{{\bm{H}}}_{l}^{-1},~~\alpha_{l}\in[0,1],$ (69)
and set
$\widehat{{\bm{W}}}_{l}^{\mathrm{QEP}}\coloneq{\bm{W}}_{l}^{\ast}(\alpha_{l})+{\bm{E}}_{l}.$ (70)
Define the per-layer pre-activation residuals
${\bm{R}}_{l}^{M}\coloneq{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}^{M}\widehat{{\bm{X}}}_{l}^{M},~~M\in\{\mathrm{BASE},\mathrm{QEP}\},$ (71)
the global Lipschitz upper bound on the output mismatch
$\mathcal{U}^{M}\coloneq\sum_{k=1}^{L}\left(\prod_{s=k+1}^{L}\gamma_{s}\|{\bm{W}}_{s}\|_{2}\right)\gamma_{k}\|{\bm{R}}_{k}^{M}\|_{F}.$ (72)
Then for every choice of $\{\alpha_{l}\}_{l=1}^{L}\subset[0,1]$,
$\mathcal{U}^{\mathrm{QEP}}\leq\mathcal{U}^{\mathrm{BASE}}.$ (73)
Consequently,
$\|{\bm{\delta}}_{L+1}^{\mathrm{QEP}}\|_{F}\leq\mathcal{U}^{\mathrm{QEP}}\leq\mathcal{U}^{\mathrm{BASE}},~~\|{\bm{\delta}}_{L+1}^{\mathrm{BASE}}\|_{F}\leq\mathcal{U}^{\mathrm{BASE}}.$ (74)
@@
Fix a method $M\in\{\mathrm{BASE},\mathrm{QEP}\}$. By Lipschitz continuity,
$\|{\bm{\delta}}_{l+1}^{M}\|_{F}=\|\sigma_{l}({\bm{W}}_{l}{\bm{X}}_{l})-\sigma_{l}(\widehat{{\bm{W}}}_{l}^{M}\widehat{{\bm{X}}}_{l}^{M})\|_{F}\leq\gamma_{l}\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}^{M}\widehat{{\bm{X}}}_{l}^{M}\|_{F}=\gamma_{l}\|{\bm{R}}_{l}^{M}\|_{F}.$ (75)
Iterating this inequality through the remaining layers yields
$\|{\bm{\delta}}_{L+1}^{M}\|_{F}\leq\sum_{k=1}^{L}\left(\prod_{s=k+1}^{L}\gamma_{s}\|{\bm{W}}_{s}\|_{2}\right)\gamma_{k}\|{\bm{R}}_{k}^{M}\|_{F}=\mathcal{U}^{M}.$ (76)
It remains to show $\mathcal{U}^{\mathrm{QEP}}\leq\mathcal{U}^{\mathrm{BASE}}$. We prove a stronger per-layer inequality:
$\|{\bm{R}}_{l}^{\mathrm{QEP}}\|_{F}\leq\|{\bm{R}}_{l}^{\mathrm{BASE}}\|_{F},~~\forall l.$ (77)
Fix $l$ and write $\widehat{{\bm{X}}}_{l}\coloneq\widehat{{\bm{X}}}_{l}^{\mathrm{QEP}}$ and ${\bm{\delta}}_{l}\coloneq{\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$. Define the orthogonal projection
${\bm{P}}_{l}\coloneq\widehat{{\bm{X}}}_{l}^{\top}(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}\widehat{{\bm{X}}}_{l},$ (78)
which satisfies ${\bm{P}}_{l}^{2}={\bm{P}}_{l}$ and ${\bm{P}}_{l}^{\top}={\bm{P}}_{l}$.
By the construction of ${\bm{W}}_{l}^{\ast}(\alpha_{l})$, we have the exact identity
${\bm{W}}_{l}^{\ast}(\alpha_{l})\widehat{{\bm{X}}}_{l}={\bm{W}}_{l}\widehat{{\bm{X}}}_{l}+\alpha_{l}{\bm{W}}_{l}{\bm{\delta}}_{l}{\bm{P}}_{l}.$ (79)
Therefore, using $\widehat{{\bm{W}}}_{l}^{\mathrm{QEP}}={\bm{W}}_{l}^{\ast}(\alpha_{l})+{\bm{E}}_{l}$,
${\bm{R}}_{l}^{\mathrm{QEP}}$ $={\bm{W}}_{l}{\bm{X}}_{l}-({\bm{W}}_{l}^{\ast}(\alpha_{l})+{\bm{E}}_{l})\widehat{{\bm{X}}}_{l}$ (80)
$={\bm{W}}_{l}(\widehat{{\bm{X}}}_{l}+{\bm{\delta}}_{l})-{\bm{W}}_{l}^{\ast}(\alpha_{l})\widehat{{\bm{X}}}_{l}-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}$ (81)
$={\bm{W}}_{l}{\bm{\delta}}_{l}-\alpha_{l}{\bm{W}}_{l}{\bm{\delta}}_{l}{\bm{P}}_{l}-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}$ (82)
$={\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha_{l}{\bm{P}}_{l})-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}.$ (83)
Consider the BASE residual at the same layer evaluated on the same $\widehat{{\bm{X}}}_{l}$:
$\widetilde{{\bm{R}}}_{l}^{\mathrm{BASE}}\coloneq{\bm{W}}_{l}{\bm{X}}_{l}-({\bm{W}}_{l}+{\bm{E}}_{l})\widehat{{\bm{X}}}_{l}={\bm{W}}_{l}{\bm{\delta}}_{l}-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}.$ (84)
Hence
${\bm{R}}_{l}^{\mathrm{QEP}}=\widetilde{{\bm{R}}}_{l}^{\mathrm{BASE}}-\alpha_{l}{\bm{W}}_{l}{\bm{\delta}}_{l}{\bm{P}}_{l}.$ (85)
Since ${\bm{P}}_{l}$ is an orthogonal projection and $0\leq\alpha_{l}\leq 1$, Lemma B.8 implies
$\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha_{l}{\bm{P}}_{l})\|_{F}\leq\|{\bm{W}}_{l}{\bm{\delta}}_{l}\|_{F}.$ (86)
Therefore, by the triangle inequality,
$\|{\bm{R}}_{l}^{\mathrm{QEP}}\|_{F}=\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha_{l}{\bm{P}}_{l})-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}\|_{F}\leq\|\widetilde{{\bm{R}}}_{l}^{\mathrm{BASE}}\|_{F}.$ (87)
Finally, since $\widehat{{\bm{X}}}_{l}^{\mathrm{BASE}}$ is the BASE activation produced by the BASE recursion, ${\bm{R}}_{l}^{\mathrm{BASE}}$ is exactly $\widetilde{{\bm{R}}}_{l}^{\mathrm{BASE}}$ evaluated at $\widehat{{\bm{X}}}_{l}^{\mathrm{BASE}}$. Thus, taking $\widehat{{\bm{X}}}_{l}=\widehat{{\bm{X}}}_{l}^{\mathrm{BASE}}$ in the above inequality yields Eq. (77). Summing with nonnegative weights $(\prod_{s=k+1}^{L}\gamma_{s}\|{\bm{W}}_{s}\|_{2})\gamma_{k}$ yields $\mathcal{U}^{\mathrm{QEP}}\leq\mathcal{U}^{\mathrm{BASE}}$, thus completing the proof. ∎
We further demonstrate that the final quantization error decreases monotonically as each propagation strength parameter $\alpha_{l}$ approaches $1$.
Corollary B.6.
Fix the layer index $l\in\{1,\dots,L\}$ and assume that $\widehat{{\bm{H}}}_{l}\coloneq\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top}$ is invertible. Let the activation mismatch be ${\bm{\delta}}_{l}\coloneq{\bm{X}}_{l}-\widehat{{\bm{X}}}_{l}$. Define the orthogonal projection
${\bm{P}}_{l}\coloneq\widehat{{\bm{X}}}_{l}^{\top}(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}\widehat{{\bm{X}}}_{l}\in{\mathbb{R}}^{m\times m}.$ (88)
For $\alpha\in[0,1]$, define the QEP corrected weight in the continuous domain as
${\bm{W}}_{l}^{\ast}(\alpha)\coloneq{\bm{W}}_{l}+\alpha{\bm{W}}_{l}{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}\widehat{{\bm{H}}}_{l}^{-1}.$ (89)
Let $\widehat{{\bm{W}}}_{l}(\alpha)\coloneq{\bm{W}}_{l}^{\ast}(\alpha)+{\bm{E}}_{l}$ be defined by a fixed matrix ${\bm{E}}_{l}$. Then, for any $0\leq\alpha^{\prime}\leq\alpha\leq 1$,
$\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha{\bm{P}}_{l})\|_{F}\leq\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha^{\prime}{\bm{P}}_{l})\|_{F}.$ (90)
Moreover, the pre-activation residual satisfies the exact identity
${\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}(\alpha)\widehat{{\bm{X}}}_{l}={\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha{\bm{P}}_{l})-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l},$ (91)
and hence the following upper bound is also monotone in $\alpha$:
$\|{\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}(\alpha)\widehat{{\bm{X}}}_{l}\|_{F}\leq\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha{\bm{P}}_{l})\|_{F}+\|{\bm{E}}_{l}\|_{2}\|\widehat{{\bm{X}}}_{l}\|_{F}.$ (92)
Proof.
First, ${\bm{P}}_{l}$ is an orthogonal projection. Indeed,
${\bm{P}}_{l}^{\top}=\widehat{{\bm{X}}}_{l}^{\top}(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}\widehat{{\bm{X}}}_{l}={\bm{P}}_{l},$ (93)
and
${\bm{P}}_{l}^{2}=\widehat{{\bm{X}}}_{l}^{\top}(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}\underbrace{\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top}}_{=\widehat{{\bm{H}}}_{l}}(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}\widehat{{\bm{X}}}_{l}={\bm{P}}_{l}.$ (94)
Thus, Lemma B.8 applies with ${\bm{Z}}\coloneq{\bm{W}}_{l}{\bm{\delta}}_{l}$ and ${\bm{P}}\coloneq{\bm{P}}_{l}$, yielding
$\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha{\bm{P}}_{l})\|_{F}\leq\|{\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha^{\prime}{\bm{P}}_{l})\|_{F},$ (95)
which proves Eq. (90).
Next, using the definitions of ${\bm{W}}_{l}^{\ast}(\alpha)$ and $\widehat{{\bm{H}}}_{l}^{-1}=(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}$,
${\bm{W}}_{l}^{\ast}(\alpha)\widehat{{\bm{X}}}_{l}={\bm{W}}_{l}\widehat{{\bm{X}}}_{l}+\alpha{\bm{W}}_{l}{\bm{\delta}}_{l}\widehat{{\bm{X}}}_{l}^{\top}(\widehat{{\bm{X}}}_{l}\widehat{{\bm{X}}}_{l}^{\top})^{-1}\widehat{{\bm{X}}}_{l}={\bm{W}}_{l}\widehat{{\bm{X}}}_{l}+\alpha{\bm{W}}_{l}{\bm{\delta}}_{l}{\bm{P}}_{l}.$ (96)
Therefore, since $\widehat{{\bm{W}}}_{l}(\alpha)={\bm{W}}_{l}^{\ast}(\alpha)+{\bm{E}}_{l}$ and ${\bm{X}}_{l}=\widehat{{\bm{X}}}_{l}+{\bm{\delta}}_{l}$,
${\bm{W}}_{l}{\bm{X}}_{l}-\widehat{{\bm{W}}}_{l}(\alpha)\widehat{{\bm{X}}}_{l}$ $={\bm{W}}_{l}(\widehat{{\bm{X}}}_{l}+{\bm{\delta}}_{l})-({\bm{W}}_{l}^{\ast}(\alpha)+{\bm{E}}_{l})\widehat{{\bm{X}}}_{l}$ (97)
$={\bm{W}}_{l}{\bm{\delta}}_{l}-\alpha{\bm{W}}_{l}{\bm{\delta}}_{l}{\bm{P}}_{l}-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}$ (98)
$={\bm{W}}_{l}{\bm{\delta}}_{l}({\bm{I}}-\alpha{\bm{P}}_{l})-{\bm{E}}_{l}\widehat{{\bm{X}}}_{l},$ (99)
which is Eq. (91). Finally, Eq. (92) follows from the triangle inequality and $\|{\bm{E}}_{l}\widehat{{\bm{X}}}_{l}\|_{F}\leq\|{\bm{E}}_{l}\|_{2}\|\widehat{{\bm{X}}}_{l}\|_{F}$. ∎
B.4 Relationship of QEP Correction and Ridge Regularization